Completing the square?
Can someone show me the steps to completing the square for this please?

- 1/5x^2 - 1/2x + 2 = 0

Found 2 solutions by ewatrrr, edjones:
Answer by ewatrrr(24785)


Hi
Completing the square:
- 1/5x^2 - 1/2x + 2 = 0 |Multiplying thru by -10 so as all denominators= 1
2x^2 + 5x -20 = 0
2(x^2 + 5/2) - 20 = 0
2[x + 5/4)^2 -25/16] - 20 = 0
2(x+5/4)^2 - 25/8 - 160/8 = 0
2(x+5/4)^2 - 185/8 = 0

Answer by edjones(8007)

- 1/5x^2 - 1/2x + 2 = 0
x^2 + 5x/2 - 10 = 0 multiply each side by -5.
x^2 + 5x/2 + 25/16 = 10 + 25/16
(x+ 5/4)^2=185/16
x + 5/4 = +-sqrt(185)/4
x=(-5+-sqrt(185))/4
x=(-5+sqrt(185))/4, x=(-5-sqrt(185))/4
.
Ed